a) The roll preparation involves two variable cost components: (i) the center filling

(ii) Ten minutes of labour

Demand curve: Q = 100 – P

Find the carnot's equilibrium


1
Expert's answer
2021-04-14T06:16:15-0400
"TR=pQ"

"TR=p(100-p)"

"TR=100p-p^2"

"MR=100-2p"

"100-2p=0"

"p=50"

"Q=50"